12. Order of operations A set of three scores consists of the values 3, 7, and 2. Σ2X-2 = 22 E(x-1)2 = -41 Hint: Remember to follow the order of mathematical operations.
Answer #1

The values are 3,7 and 2.

Σ2X - 2

= 2ΣX - 2

= 2 * (3 + 7 + 2) - 2

= 2 * 12 - 2

= 24 - 2

= 22.

Σ(X - 1)2

= Σ(X2 - 2X + 1)

= ΣX2 - Σ2X + Σ1

= ΣX2 - 2ΣX + Σ1

= (32 + 72 + 22) - 2(3 + 7 + 2) + 3

= (9 + 49 + 4) - 2 * 12 + 3

= 62 - 24 + 3

= 38 + 3

= 41.